22.30 was spent on a total of 14 tickets,if adults was 2.25, children admission was1.10.How many children attended the movie.
Answer by Maths68(1474)

Let
x=Adult Ticket
y=Children Ticket
Cost of Adult ticket = 2.25
Cost of Children ticket = 1.10
Cost of x tickets sold=2.25*x
Cost of y tickets sold=1.10*y
Adult tickets sold + children’s tickets sold = Total tickets sold
x+y=14
x=14-y………………(1)
Cost of Adult tickets sold + Cost of Children’s tickets sold = Cost of total tickets sold
2.25x+1.10y=22.30……..(2)
Put value of x from (1) to (2)
2.25x+1.10y=22.30……..(2)
2.25(14-y)+1.10y=22.30
2.25(14-y)+1.10y=22.30
31.5-2.25y+1.10y=22.30
-1.15y=22.30-31.5
-1.15y=-9.2
y=-9.2/-1.15
y=8
8 children attended the movie.